OpenStudy (anonymous):

The length of a pool must be 3x the width plus 3. The depth must be 7 less than 2x the length. How do I write the depth in an algebraic expression? x=width 3x+3=length How do I write the depth? I got 2(3x+3)-7 -- is this correct?

OpenStudy (nirmalnema):

yes you are correct

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

Thank you both!

OpenStudy (nirmalnema):

welcome

OpenStudy (anonymous):

How do I write the volume? I know the formula for the volume of a rectangular prism is LengthxWidthxHight(or depth)

OpenStudy (anonymous):

Volume = length x width x height

OpenStudy (nirmalnema):

(x)*(3x+3)*(2(3x+3)-7)

OpenStudy (anonymous):

Volume = x (3x+3) 2(3x+3) - 7 = 2x(3x+3)(3x+3) - 7

OpenStudy (anonymous):

Thank you @Easyaspi314 !!!

OpenStudy (anonymous):

and @nirmalnema !! you guys helped me a lot (:

OpenStudy (phi):

I got 2(3x+3)-7 -- is this correct? yes, that is the depth you could simplify it by distributing the 2 2(3x+3)-7 6x + 6 -7 6x-1 <-- another way to write the depth, after simplifying

OpenStudy (phi):

with width = x , length= 3x+3= 3(x+1) and depth= (6x-1) the volume is width * length * depth x * 3(x+1)(6x-1) or 3x(x+1)(6x-1)

OpenStudy (anonymous):

@phi I don't under stand how the depth=(6x-1) I understand the 6x but why is it -1 when its supposed to be -7?

OpenStudy (anonymous):

@phi nevermind! I didn't see your first reply!! Thank you guys so much!

OpenStudy (anonymous):

Now I have to classify the volume polynomial by degree and number of terms. 3x(x+1)(6x-1) (3x^2+3x) (6x-1) 18x^3-3x^2+18x^2-3x 18x^3+15x^2-3x 3rd degree trinomial. Is this right? I'm not so confident about this one..

OpenStudy (anonymous):

@phi @Easyaspi314 @nirmalnema

OpenStudy (nirmalnema):

yup you r right..

OpenStudy (anonymous):

Thanks again! lol

OpenStudy (anonymous):

Bernard is realizing that his dimensions are not fitting the way that he wants it to because he forgot about the deck. The deck is going to be a rectangular space that borders the pool. Also, for the pool to fit just right there must be a special product represented when either the area or volume expressions are determined. He wants you to design your own rectangular pool using polynomial expressions, including the deck to borders the pool. a.Include a diagram or sketch of your pool and deck. Your diagram must have all dimensions represented with polynomials and ensure that a special product will be used. How do I use a special product for this? @Easyaspi314 @nirmalnema